3. Both the flammability reaction and the explosion reaction involve the combination of acetylene with oxygen. Write the correct equation for each. (a) Which of these represents INCOMPLETE COMBUSTION? Upon what evidence do you make this choice? C242 C2 on The equation is provided. I just need to know the second part of the question “Upon what evidence do you make this choice”. Please explain clearly and with facts.
Answer #1

Incomplete combustion is typically accompanied with CO(cabron monoxide) production

C2H2 + O2 = CO+ H2O

balance C, H, then O

C2H2 + 3/2O2 = 2CO+ H2O

then...

we can ensure this is NOT 100% complete, since CO will burn with oxygen if present

CO(g) + 1/2O2(g) = CO2(g)

if this is 100% oxidized/reacted, then you should NOT see any flame formation, if it has CO , then a flame will appear